Wow, it’s just like Example 2 all over again! In Problem 30, you’re given the equation:
f(x) = 4x – 3
for which you need to solve the following:
30) f(-1)

Answer by SandySharma(35)

f(x) = 4x – 3
Find f(-1)
Replace x = -1 in given function
f(-1) = 4(-1) – 3 = -4 - 3= -7
f(-1) = -7